40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Mar 20 12:42:01 2015, Jenne, Update, IOO, Waking up the IFO 
    Reply  Fri Mar 20 13:29:33 2015, Koji, Update, IOO, Waking up the IFO 
    Reply  Fri Mar 20 16:44:49 2015, ericq, Update, IOO, Waking up the IFO 
       Reply  Fri Mar 20 23:37:46 2015, Jenne, Update, SUS, Waking up the IFO 
          Reply  Sat Mar 21 05:19:49 2015, Jenne, Update, LSC, IFO awake carm_cm_up_zip.sh.gz
          Reply  Sun Mar 22 13:53:44 2015, rana, Update, SUS, ITMX alignment jumps 
Message ID: 11154     Entry time: Sat Mar 21 05:19:49 2015     In reply to: 11153
Author: Jenne 
Type: Update 
Category: LSC 
Subject: IFO awake 

[Jenne, Den]

The problem with the ASS turned out to be a mode that was rung up at 1326Hz in ETMX.  It was rung up when the Xarm's overall gain was too high.  So, by turning down the digital gain we were able to prevent it ringing up, and then the ASS worked.  To circumvent this, we added a notch to the violin filter bank.  It turned out that, upon trying to check if this existed also for the Yarm by turning up the digital gain, the ETMY frequency was almost identical.  So, the same single notch is in both ETMs, and it covers the modes for both ETMs.

After that, we got back to locking.  We have made at least 9 transitions to all-RF (both CARM and DARM) tonight (I have lost track of how many Den has done while I've been writing this - maybe we're up to 10 or so.). We have changed the order of things a little bit, but they're mostly similar to last week.  There are some new notches in the CARM_B filter bank, as well as a 700Hz low pass.  We have not been using the lead filter in DARM from last week.  Script is checked in, and also zipped and attached.  At first CARM was actuating on ETMs, but the last half of the locks we've been using MC2.  The script is optimized for MC2 actuation.

While locked all RF, we phased REFL55 in preparation for transitioning PRMI over from REFL165. REFL55 phase was +125, now is +80, give or take 5 deg.  We have tried measuring the relative gain and sign between REFL55 and REFL165, but we keep losing lock, perhaps as a result of the TFs Den is taking.  He's being gentle though. 

Up next:

Transition PRMI

Measure CARM loop (why was SRmeasure not working?? is it plugged in??)

Turn on AO boosts, etc.

 

Attachment 1: carm_cm_up_zip.sh.gz  1 kB
ELOG V3.1.3-